OpenStudy (anonymous):

Find the coordinates of the point that lies one fourth of the way from (-10,6) to (0,-2)

OpenStudy (amistre64):

find the length (L) and use it as a ruler

OpenStudy (amistre64):

-10 6 0 2 ------ -10 8 100+64 .... lngth = sqrt(164)

OpenStudy (amistre64):

hmmm, i wonder a full slope is -8/10 if we divide those by 4 we should get 1/4 of the slope -8/4 = -2 10/4 = 5/2 use the staring point and add this modified slope to it -10 6 5/2 -2 ----------- -15/2 4 maybe